Part of $12000 is invested in 9% bonds and the remainder in common stocks paying 8%. if the yearly income from both investments is $1,04.00, how much was invested in bonds and how much was invested in stocks?
Answer by richwmiller(17219)

x+y=12000,
0.08*x+0.09*y=1040
x=12000-y
0.08*(12000-y)+0.09*y=1040
960-0.08y+0.09*y=1040
0.01*y=80
y=8000 at 9%
x=12000-y
x=4000 at 8%
check
0.08*4000+0.09*8000=1040
320+720=1040
1040=1040
ok